OpenStudy (anonymous):

HELP ILL DO ANYTHING - - ---- Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). (2 points) y = 1 over 3 x + 7 y = -3x + 3 y = 1 over 3 x + 3 y = -3x + 7

OpenStudy (anonymous):

\[y-6=3(x+1)\] is a start

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

solve for \(y\) and you are done

OpenStudy (anonymous):

@satellite73 i need your help still im stupid

OpenStudy (anonymous):

@satellite73 pleaseeeeeee im in a rush

OpenStudy (anonymous):

help :(

OpenStudy (anonymous):

Since the slope of parallel lines has to be the same you can do y=-3x+b. Then plug in the point you are given and solve for b to find the y intecept

OpenStudy (anonymous):

ok can you help with that because im really dumb

OpenStudy (anonymous):

and explain

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

y=-3x+b plug in the give point (-1,6). Since ordered pairs are in (x,y) form -1=x and 6=y 6=-3(-1)+b Then solve for b 6=3+b subtract 3 from both sides and you get 3=b. Since the slope of the parallel line is the same your parallel equation is y=-3x+3

OpenStudy (anonymous):

i apologize but im in a rush :(

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

can you help me with a few more?

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 3 over 2x + 4 and passes through the point (-6, 3). (2 points) y = 3 over 2x - 1 y = 3 over 2x + 3 y = - 2 over 3x - 1 y = - 2 over 3x + 3

OpenStudy (anonymous):

can you guide me step by step?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

yeah just a sec

OpenStudy (anonymous):

ok if i dont answer that means i need a sec need to get a beverage'

OpenStudy (anonymous):

sorry about that im back!

OpenStudy (anonymous):

Is this your equation  or is it this